Total Manufacturing has an expected EBIT of $40,000 per year in perpetuity and a tax rate...

  1. Total Manufacturing has an expected EBIT of $40,000 per year in perpetuity and a tax rate of 20%. The firm currently has no debt. Its cost of debt is 8% and unlevered cost of capital is 14%. If the firm changes its capital structure by borrowing $120,000 to repurchase the same amount of equity, what would be the firm's value under the new capital structure?
Answer #1

Solution :-

Value of Unlevered Firm is calculated = EBIT * ( 1 - Tax ) / UL Ke   

Where EBIT = $40,000

Tax Rate = 20%

UL Ke ( Unlevered Cost of Capital ) = 14%

Now Value of Unlevered Firm = $40,000 * ( 1 - 0.20 ) / 0.14 = 228,571.40

Value of Levered Firm = Value of Unlevered Firm + ( Debt * Tax )

Value of Levered Firm = 228,571.40 + ( $120,000 * 0.20 ) = $252,571.4

Firm Value = $252,571.40